OpenStudy (anonymous):

I ONLY HAVE 1 HOUR. PLEASE HELP! MEDAL!!! What is the slope of the line represented by the equation 4x + 8y = 4 ?

OpenStudy (anonymous):

The slope is -1/2.

OpenStudy (anonymous):

THANK YOU!! Do you have time for more?

OpenStudy (anonymous):

shoot 'em (the questions)

OpenStudy (anonymous):

What is the simplified form of the expression? 20y – 2(3y – 10)

OpenStudy (anonymous):

14y+20

OpenStudy (anonymous):

THANK YOU!!!! What is the slope of the line represented by the equation Y= x/6 -5

OpenStudy (anonymous):

Y=mx+b where m is the slope so m = 1/6

OpenStudy (anonymous):

You are seriously a LIFE SAVER!!!!

OpenStudy (anonymous):

How many solutions does this system of equations have? y = –2x + 3 y = –2x – 1 A)0 B)1 C) 2 D) infinitely many

OpenStudy (mathmale):

Lalo: You are an answer-man, all right. Pleae consider guiding Tati through the steps of the problem solutions rather than just providing her with the answers.

OpenStudy (anonymous):

If you substitute one equation on the other one you can only get one answer

OpenStudy (mathmale):

My point is that you are ahead of Tati in your knowledge of algebra. So what may look easy for you may be not at all clear for Tati. That's why I suggest you lead her through solving problems, instead of providing her with answers. Surely appreciate your involvement here. Thank you.